OpenStudy (i_love_my_nieces):

When baking a cake, you have a choice of the following pans: a round cake pan that is 2 inches deep and has a 7 inch diameter. a 6 inch x 9 inch rectangular cake pan that is 2 inches deep a. Which of these pans has the larger volume? Justify your answer. b. How much icing will be needed to frost each cake, excluding the bottom of the cake? Justify your answer.

OpenStudy (anonymous):

i bet you guys can walk her through.. :P

OpenStudy (austinl):

How do you find the volume of a cylinder? How do you find the volume of a rectangular prism?

OpenStudy (i_love_my_nieces):

lwh

OpenStudy (austinl):

\(V_{rec}=l\times w\times h\) What is a cylinder?

OpenStudy (austinl):

\(V_{cyl}=\pi r^2 h\)

OpenStudy (austinl):

Find the volume of each of them and see which one is larger.

OpenStudy (i_love_my_nieces):

Thank you @austinL I got the answer!! ;) 0.o

OpenStudy (austinl):

did you get the answer to part b?

OpenStudy (i_love_my_nieces):

Yepper-doodle-doo.

OpenStudy (austinl):

Good job :)

OpenStudy (i_love_my_nieces):

Thank you, can you help me with 2 more?